# Spin-2 fields and cosmic censorship

### v1

@@ VI The correct application of Sorce-Wald conditions
The derivation of (51) is also given in the Appendix. The expression on the right-hand-side of (51) is identical with the self-energy derived from Will’s argument will based on the induced increase in the angular velocity of the event horizon.
@@
We previously used these backreaction effects in the overspinning problems involving bosonic and fermionic fields kerrmog ; spinhalf ; threehalves . The fact that the two results coincide lends credence to the validity of the methods to derive the backreaction effects.
@@ VII Conclusions
We analysed the case of nearly extremal black holes defining the closeness to extremality to be second order in the small parameter $\epsilon$. We derived a range of frequencies that could overspin the nearly extremal black hole if the absorption probability was of the order of unity. This is a narrow range of width $\sim\epsilon$ bounded below by $\omega=m\Omega$ and above by $\omega=(1/M)$. We showed that the maximum value of the absorption probability for these challenging modes is of the order $\epsilon^{5}$. A very small fraction of the test field is absorbed by the black hole. Therefore the contribution to the mass and the angular momentum parameters is very small. It is not possible to surpass the gap of width $\sim\epsilon^{2}$ to drive the black hole to extremality or beyond. When one incorporates the absorption probabilities, the cosmic censorship conjecture remains valid without the need to employ the back-reaction effects.
@@
We also reviewed and extended our previous comments on Sorce-Wald method. Previously we stated that the function $f(\lambda)$ defined by Sorce and Wald involves order of magnitude problems, as one is forced to multiply the contribution of the second order perturbations by the square of an extra non-physical parameter $\lambda$ absorp . Here, we pointed out that the conditions derived by Sorce and Wald correctly describe the interaction of black holes with test particles and fields. The first condition (25) was first derived by Needham in 1980 needham (and independently by different authors, see e.g. natario ). The condition on the second order perturbations (26) is an original contribution by Sorce and Wald. The backreactions effects can be derived by using the Sorce-Wald condition and incorporated into the analysis. To remedy the order of magnitude problems, one simply abandons $f(\lambda)$ and follows the line of research developed by Semiz and Düztaş. We applied this method to the problems of overcharging Reissner-Nordström black holes and overspinning Kerr black holes. First we derived the Sorce-Wald condition for Reissner-Nordström and Kerr black holes. It turns out that the condition for Kerr black holes is identical with the self-energy derived from Will’s argument will , which we have used in our previous works. We showed that overcharging and overspinning is possible when one ignores backreaction effects, and backreaction effects derived from the Sorce-Wald condition fixes the problem, in accord with previous results derived by Hubeny hu , Jacobson-Sotiriou js , and Düztaş-Semiz overspin . Overspinning and overcharging occurs in the range $\delta M\sim\epsilon$, for which $f(\lambda)$ defined by Sorce and Wald acquires an invariant value of $M^{2}\epsilon^{2}$ and conveys no information about the interaction of black holes with test particles and fields. The legitimate way to make use of the Sorce-Wald conditions is to abandon $f(\lambda)$ and directly incorporate the backreactions into the analysis.
@@
Needham’s condition gives the lower bound for the energy of the perturbation to allow its absorption by the black hole. However, one cannot directly infer that the energy of the optimal perturbations should have the value of the lower bound. In a more subtle approach, one discerns that the perturbations with the lowest energy are entirely reflected back to infinity since their absorption probability is zero. For a test field, Needham’s condition implies that the lower bound for the frequency is $\omega=m\Omega$. In the classical picture this field is entirely reflected back to infinity. Therefore the absorption probability should be taken into consideration to identify the optimal perturbations. In the examples using Needham and Sorce-Wald conditions in section (VI), we ignored the absorption probabilities. In that case the derivation for the overspinning problem would be identical for test bodies, test particles, and test fields with spin-0, spin-1, and spin-2. A more accurate derivation should involve the explicit form of the absorption probabilities as executed in sections (III) and (IV). The problem is that the absorption probability is not well-defined for test bodies and the probabilities for test fields do not apply to test particles. (The latter is a subtle point which we elucidated in q7 .) The validity of the results ignoring the absorption probabilities is restricted to test bodies for which the absorption probability appears least relevant.
@@
Finally we would like to point out that the results for the perturbations that do and do not satisfy the null energy condition should not be confused. For perturbations satisfying the null energy condition, there exists a lower bound for the energy to allow the absorption of the test particle or the field by the black hole. The lower bound can be derived from Needham’s condition (25) which pre-assumes that the null energy condition is satisfied. Fermionic fields do not satisfy the null energy condition and there exists no lower bound for the energy of the perturbation to allow its absorption by the black hole. An equivalent statement is that superradiance does not occur for fermionic fields. To be more precise the absorption probability is always positive, as justified by Page’s results page . The contribution of the lower energy modes to the angular momentum or charge parameters is much larger than their contribution to the mass parameter. The absorption of these modes by the black hole leads to a generic violation of the cosmic censorship conjecture which cannot be fixed by backreaction effects duztas ; toth ; generic ; spinhalf ; threehalves . This does not contradict the fact that cosmic censorship remains valid for perturbations satisfying the null energy condition, which can be considered complete with the results for spin-2 fields derived in this work.
Appendix
@@
In this section we evaluate the Sorce-Wald condition (26) for a Reissner-Nordström and a Kerr black hole parametrized as (27). The surface gravity is given by (see Equation (116) in w2 ):
@@
The second order variation in the area of the black hole is given by (see Equation (113) in w2 ):

### v2

@@ VI The correct application of Sorce-Wald conditions
The derivation of (52) is also given in the Appendix. The expression on the right-hand-side of (52) is identical with the self-energy derived from Will’s argument [68] based on the induced increase in the angular velocity of the event horizon.
@@
We previously used these backreaction effects in the overspinning problems involving bosonic and fermionic fields [59, 63, 64]. The fact that the two results coincide lends credence to the validity of the methods to derive the backreaction effects.
@@ VII Conclusions
We analysed the case of nearly extremal black holes defining the closeness to extremality to be second order in the small parameter $\epsilon$. We derived a range of frequencies that could overspin the nearly extremal black hole if the absorption probability was of the order of unity. This is a narrow range of width $\sim\epsilon$ bounded below by $\omega=m\Omega$ and above by $\omega=(1/M)$. We mentioned that the highest absorption probability for $m=2$ modes pertains to spin-2 fields. In that respect one may expect spin-2 fields to be a better candidate for the possibility to destroy the event horizon. However, we showed that the maximum value of the absorption probability for these challenging modes is of the order $\epsilon^{5}$. A very small fraction of the test field is absorbed by the black hole. Therefore the contribution to the mass and the angular momentum parameters is very small. It is not possible to surpass the gap of width $\sim\epsilon^{2}$ to drive the black hole to extremality or beyond. When one incorporates the absorption probabilities, the cosmic censorship conjecture remains valid without the need to employ the back-reaction effects.
The absorption probability for $m=2$ modes for spin-2 fields is even higher than hypothetical spin-3 fields since the major contribution comes from $l=s$ modes. For the same reason the highest absorption probability for $m=3$ modes pertains to spin-3 fields. The natural question here is how we can extrapolate our results for the spin-2 case to hypothetical higher integer spin fields. In particular for spin-3 fields the absorption probability for $m=3$ modes will involve a term like $(\omega-m\Omega)^{7}$, therefore it will be of the order $\epsilon^{7}$. Though this is the highest probability for $m=3$ modes, it practically implies that $m=3$ modes will not be absorbed by the black hole. Therefore the results derived in this work imply that spin-3 and higher integer spin fields do not challenge the validity of cosmic censorship.
@@
We also reviewed and extended our previous comments on Sorce-Wald method. Previously we stated that the function $f(\lambda)$ defined by Sorce and Wald involves order of magnitude problems, as one is forced to multiply the contribution of the second order perturbations by the square of an extra non-physical parameter $\lambda$ [67]. Here, we pointed out that the conditions derived by Sorce and Wald correctly describe the interaction of black holes with test particles and fields. The first condition (26) was first derived by Needham in 1980 [69] (and independently by different authors, see e.g. [31]). The condition on the second order perturbations (27) is an original contribution by Sorce and Wald. The backreactions effects can be derived by using the Sorce-Wald condition and incorporated into the analysis. To remedy the order of magnitude problems, one simply abandons $f(\lambda)$ and follows the line of research developed by Semiz and Düztaş. We applied this method to the problems of overcharging Reissner-Nordström black holes and overspinning Kerr black holes. First we derived the Sorce-Wald condition for Reissner-Nordström and Kerr black holes. It turns out that the condition for Kerr black holes is identical with the self-energy derived from Will’s argument [68], which we have used in our previous works. We showed that overcharging and overspinning is possible when one ignores backreaction effects, and backreaction effects derived from the Sorce-Wald condition fixes the problem, in accord with previous results derived by Hubeny [4], Jacobson-Sotiriou [5], and Düztaş-Semiz [29]. Overspinning and overcharging occurs in the range $\delta M\sim\epsilon$, for which $f(\lambda)$ defined by Sorce and Wald acquires an invariant value of $M^{2}\epsilon^{2}$ and conveys no information about the interaction of black holes with test particles and fields. The legitimate way to make use of the Sorce-Wald conditions is to abandon $f(\lambda)$ and directly incorporate the backreactions into the analysis.
@@
Needham’s condition gives the lower bound for the energy of the perturbation to allow its absorption by the black hole. However, one cannot directly infer that the energy of the optimal perturbations should have the value of the lower bound. In a more subtle approach, one discerns that the perturbations with the lowest energy are entirely reflected back to infinity since their absorption probability is zero. For a test field, Needham’s condition implies that the lower bound for the frequency is $\omega=m\Omega$. In the classical picture this field is entirely reflected back to infinity. Therefore the absorption probability should be taken into consideration to identify the optimal perturbations. In the examples using Needham and Sorce-Wald conditions in section (VI), we ignored the absorption probabilities. In that case the derivation for the overspinning problem would be identical for test bodies, test particles, and test fields with spin-0, spin-1, and spin-2. A more accurate derivation should involve the explicit form of the absorption probabilities as executed in sections (III) and (IV). The problem is that the absorption probability is not well-defined for test bodies and the probabilities for test fields do not apply to test particles. (The latter is a subtle point which we elucidated in [28].) The validity of the results ignoring the absorption probabilities is restricted to test bodies for which the absorption probability appears least relevant.
@@
Finally we would like to point out that the results for the perturbations that do and do not satisfy the null energy condition should not be confused. For perturbations satisfying the null energy condition, there exists a lower bound for the energy to allow the absorption of the test particle or the field by the black hole. The lower bound can be derived from Needham’s condition (26) which pre-assumes that the null energy condition is satisfied. Fermionic fields do not satisfy the null energy condition and there exists no lower bound for the energy of the perturbation to allow its absorption by the black hole. An equivalent statement is that superradiance does not occur for fermionic fields. To be more precise the absorption probability is always positive, as justified by Page’s results [65]. The contribution of the lower energy modes to the angular momentum or charge parameters is much larger than their contribution to the mass parameter. The absorption of these modes by the black hole leads to a generic violation of the cosmic censorship conjecture which cannot be fixed by backreaction effects [60, 61, 62, 63, 64]. This does not contradict the fact that cosmic censorship remains valid for perturbations satisfying the null energy condition, which can be considered complete with the results for spin-2 fields derived in this work.
Appendix: Sorce-Wald condition for Reissner-Nordström and Kerr black holes
@@
In this section we evaluate the Sorce-Wald condition (27) for a Reissner-Nordström and a Kerr black hole parametrized as (28). The surface gravity is given by (see Equation (116) in [66]):
@@
The second order variation in the area of the black hole is given by (see Equation (113) in [66]):